# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v6

@@ 3.4 Memorization
To prevent models from using brute-force memorization, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints—both in time and memory. In AGITB, a task typically involves predicting a temporal sequence of $n$ binary inputs, each consisting of $k$ bits. This defines a raw combinatorial space of size $|S|=2^{nk}$, representing all possible binary input sequences.
@@
AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps. This constraint significantly narrows the set of valid input sequences. The number of distinct length-$n$ temporal input sequences of $k$ bits, with the rule that a 1 never survives to the next time step, is $|S\textquoteright|=(F_{n+2})^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
@@
AGITB further reduces this space by requiring sequences to be cyclic, meaning that the first input must also satisfy the refractory constraint relative to the last input in the sequence. The number of distinct cyclic temporal input sequences with refractory constraint is given by $|S^{\prime\prime}|=(L_{n})^{k}$, where $L_{i}$ denotes the $i$-th Lucas number, with $L_{0}=2$.
Consider an illustrative example involving the learning of a temporal sequence consisting of $n=5$ input steps, each represented by a 10-bit binary vector. The total number of possible sequences is $|S|=2^{50}$. Introducing the refractory constraint significantly reduces this space to $|S\textquoteright|=(F_{7})^{10}=13^{10}$ valid sequences. Adding the cyclic condition further restricts the space to $|S^{\prime\prime}|=(L_{5})^{10}=11^{10}$ sequences.
Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality. Given that real-world sensory inputs can span tens of thousands of bits, brute-force memorization strategies become computationally intractable. This exponential complexity underscores the need for models that employ generalizable, pattern-based learning mechanisms rather than rote memorization.
@@ 3.5 Performance
Before assessing AGITB’s utility, it is important to recognize its role as a pragmatic benchmark, not as an end goal in itself. Like the Turing Test, which Harnad (1992) described as an empirical criterion rather than a philosophical sleight of hand, AGITB is intended to serve as a practical tool for evaluating progress toward AGI. The ultimate objective remains AGI’s development, not merely passing the test.
AGITB can yield meaningful insights only when developers adhere faithfully to its core requirements. Misinterpreting elements such as the “empty” initial state can lead to unreliable or misleading results.
AGITB provides a structured framework for empirically evaluating the AGI potential of various systems, including classical symbolic architectures, artificial neural networks (ANNs), and large language models. However, before benchmarking artificial systems, it is first necessary to consider the baseline performance of human cognition.
@@ 3.5.1 Human Performance
The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans. However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most requirements. Demand 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
Due to prior experience and cognitive bias, an adult human may appear to “fail” the first prerequisite (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction. In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
AGITB includes the refractory period requirement (6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution. Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection. Temporal flexibility (Requirement 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations. The property of stagnation (Requirement 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge. The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event. Because humans learn different topics at varying speeds, the rationale for the content sensitivity (Requirement 9) and the context sensitivity (Requirement 10) is self-evident. The inability to directly inspect human mental states is captured in the unobservability requirement (Requirement 11). The generalization test (Requirement 12) requires what humans naturally excel at: generalizing when faced with new but distorted input. Finally, bounded prediction latency (Requirement 13) is a hallmark of the human brain.
@@ 3.5.2 Classical Symbolic Programs
AGI approaches rooted in classical programming are fundamentally constrained by the Symbol Grounding Problem (SGP), as explained by Harnad (1990). While such systems may exhibit weak AI capabilities—producing behavior that appears intelligent—they lack grounded understanding and, therefore, cannot be considered genuine instances of AGI, as argued by Searle (1980).
In classical AI systems, a model’s initial state is inseparable from its program, which encodes fixed logic and predetermined responses. Consequently, such systems cannot begin from an unbiased or knowledge-free state, as their behavior is entirely dictated by human-authored instructions. AGITB’s first test, however, introduces a paradox: the system must initialize without any such program—that is, it must begin in a truly neutral state, free from prior knowledge or assumptions. If they adhere to the empty-state requirement, they remain inert; if they do not, they reintroduce the symbol grounding problem. In the language of Searle’s Chinese Room Argument, the ledger must be blank—containing no predefined instructions for interpreting Chinese symbols. Yet classical systems require a program to function; absent such instructions, they halt.
AGITB’s initial test formalizes the requirement to learn from scratch: it metaphorically demands that Searle acquire a language—such as Chinese—purely through exposure to input, without access to any prior rules. To enforce this condition reliably, the model comparison must be rigorous, ensuring that all relevant internal properties of the system are examined. Any structural discrepancy or hidden initialization could conceal bias and compromise the integrity of the test.
It is essential that models begin from a genuinely unbiased state. As an external framework, AGITB cannot independently verify whether a program that declares itself “empty” is truly devoid of prior knowledge. Any attempt to obscure bias at initialization compromises the integrity of the benchmark and invalidates the results of subsequent tests. By embedding prior assumptions, the developer effectively undermines the model’s claim to general, input-driven learning.
@@ 3.5.3 Artificial Neural Network Performance
Unlike classical symbolic programs, artificial neural networks do not encounter the same initialization paradox. Perceptron-based architectures are inherently robust: they always remain operational, as neurons continuously perform computations, even when their input remains neutral (for example, when an activation function returns zero).
However, the requirement that a model begins in a completely unbiased, knowledge-free state stands in fundamental tension with the prevailing paradigm of modern deep learning. Contemporary approaches typically rely on pretraining, which adjusts the network’s weights in advance through exposure to structured or labeled data. What remains absent is a mechanism by which an uninitialized network can autonomously begin learning—without supervision, without predefined semantics, and without relying on external scaffolding.
Although an untrained network may technically operate without internal excitation, no learning dynamics are triggered. The neurons are not halted but dormant and passively awaiting informative input. This reveals a critical limitation of current ANN designs: they lack the intrinsic capability to initiate unbiased, autonomous learning from scratch in arbitrary, unfamiliar environments.
@@ 3.5.4 Large Language Model Performance
Because an LLM is essentially an ANN, it fails the genesis test. This failure immediately indicates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained. Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—any such evaluations are moot until the first requirement is satisfied.
If LLMs do not qualify as AGI, could they nevertheless simulate one? For example, what if an LLM were prompted to imitate a cortex that receives binary inputs, as specified by AGITB? A structured example of such a prompt is provided in the Appendix. However, before evaluating this type of simulation, we must revisit the sensitivity (Requirement 4), which stipulates that a model must remain sensitive to any single input, even after processing an arbitrarily long sequence of subsequent inputs.
This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window. Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to forget or attenuate the long-term dependencies essential to satisfying the test.
More broadly, LLMs cannot acquire new knowledge—or learn a new language—purely from prompts. Nor can they develop the kind of grounded, context-dependent understanding of the world that humans possess. These limitations also extend to large reasoning models (LRMs), which are built upon the same underlying architecture.
@@ 3.6 Remarks
AGITB evaluates a model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs. Random input sequences with arbitrary internal correlations are used to minimize reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
This low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador et al. (2023), where cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT. AGITB’s “all-tests-must-pass” standard enforces a rigorous, biologically motivated foundation for evaluating AGI, grounded in the same principles that underlie natural intelligence.
While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework. As a recently proposed benchmark, AGITB has not yet been passed by any symbolic system, and it remains an open question whether such a solution is even theoretically possible. If a purely symbolic system were to succeed, it would suggest that AGITB fails to effectively differentiate narrow AI from genuinely general intelligence, given that symbolic systems—while proficient in syntactic manipulation—arguably lack true semantic understanding, as discussed by Searle (1980). Until such a system is demonstrated, AGITB’s assertions may be regarded as necessary (though not sufficient) conditions for identifying genuine, domain-independent intelligence. Although passing AGITB does not equate to achieving full AGI—since it does not, for instance, assess social reasoning or natural language fluency—failure to meet its conditions likely indicates a deficiency in foundational general cognitive capabilities.
Hand-engineered, task-specific systems have historically struggled to scale toward true generalization, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks. However, state-of-the-art networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input. This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB framework?
@@ 3.7 Cheating the Benchmark
Since AGITB tests are not inherently difficult to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific models and deploy them selectively based on the task at hand.
@@
While it is theoretically possible to infer a model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research. Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the Cortex programming interface—these measures would not only deter cheating but also risk making tests less intelligible to human developers. This would undermine two of the testbed’s core objectives: transparency and interpretability.
@@
For this reason, the AGITB reference implementation remains deliberately readable and accessible. To date, no classical symbolic AI system has demonstrated the flexibility or generality required to handle the arbitrary patterns of structure and length that AGITB demands. Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing an obfuscated version of the test.
4 AGITB and ARC: A Comparison
The absence of a unified definition of intelligence has led to AGI benchmarks that often suffer from conceptual ambiguity and limited practical utility. Most offer little insight into the mechanisms underlying generalization. In this regard, AGITB takes a distinct approach. Focusing on the emergence of intelligence from low-level signal-processing building blocks offers a biologically grounded and mechanistically interpretable framework.
@@ 4 AGITB and ARC: A Comparison
Among competing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization. ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recoloring, rearranging, or modifying spatial patterns in a spatial grid.
@@
However, ARC implicitly assumes the presence of high-level cognitive priors, such as object permanence, spatial reasoning, numerical abstraction, and causal inference. These priors are left undefined, placing an ambiguous burden on the model architecture. In contrast, AGITB adopts a fundamentally different perspective: it treats the neocortex—the primary focus of the benchmark—as a blank slate that must develop structure and function exclusively through interaction with temporally structured input. Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical “old brain,” as described by (Hawkins and Blakeslee, 2004, p. 66).
@@
While ARC presumes temporal reasoning, it does not explicitly test it, as it uses only two images to demonstrate a transformation. In contrast, AGITB evaluates cognition as a dynamic, unfolding process over time. A model can only develop knowledge and predictive capabilities through time, not from disconnected before-and-after snapshots that lack the temporal continuity needed to infer causal structure. For instance, to learn about objects and to recognize the object transformation, such as “move left,” a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything. Learning requires exposure to temporally structured data, enabling the model to discover invariants and causal relations through experience.
@@
In summary, ARC presupposes intelligence, whereas AGITB requires it to emerge from first principles. ARC evaluates performance based on behaviors grounded in human-like cognitive priors, while AGITB measures a system’s ability to develop necessary priors autonomously through exposure to raw input. Benchmarks such as ARC and the Turing test implicitly frame intelligence through the lens of human cognition, thereby embedding anthropocentric biases into their evaluation criteria. In contrast, AGITB conceptualizes intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions. Table 1 outlines the key differences between the two benchmarks.
Property ARC AGITB
Learning mode Online Online
Susceptible to memorization No No
Interface modality Visual Binary
Target intelligence type Human Universal
Assumed cognitive priors Yes No
Model development feedback No Yes
Fully automatable Limited Yes
Task type Synthetic Synthetic
Task production Manual Automatic
Task inspiration High-level reasoning Low-level processing
Task spatial dimensions 22 $\infty$
Task temporal structure No Yes
Table 1: Comparison of core properties between the ARC and AGITB benchmarks.

### v7

@@ 3.4 Search space
To prevent models from relying on brute-force memorisation, a robust AGI benchmark must define a problem space large enough to exceed the capacity of any model operating under realistic computational constraints in both time and memory. In AGITB, each task typically involves predicting a random temporal sequence of seven binary inputs, each comprising ten bits. This defines a raw combinatorial space of size $|S|=2^{70}$, representing all possible binary input sequences.
@@
The refractory period is a biologically inspired constraint that prohibits any neuron (bit) from firing in consecutive time steps. This restriction substantially reduces the number of valid input sequences. There are $|S^{\prime}|=(F_{7+2})^{10}=34^{10}\approx 2^{50}$ distinct seven-step temporal sequences of ten bits, assuming that a $1$ never carries over to the next time step, where $F_{i}$ is the $i$-th Fibonacci number with $F_{0}=0$.
@@
In some cases, AGITB further constrains the space by requiring the sequence to be cyclic, such that the first input also satisfies the refractory condition relative to the last input in the sequence. The number of distinct cyclic temporal input sequences with the refractory constraint is $|S"|=(L_{7})^{10}=29^{10}\approx 2^{48}$, where $L_{i}$ denotes the $i$-th Lucas number with $L_{0}=2$.
The use of length-seven sequences with ten-bit inputs is considered sufficient to detect non-AGI behaviour while maintaining computational efficiency. Increasing these default values could exceed the capabilities of a first-generation AGI under evaluation, potentially producing false negatives and substantially increasing execution time. The current configuration ensures that each test remains computationally feasible and diagnostically informative.
Within the comparatively constrained AGITB environment, every randomly generated input sequence is, in principle, learnable through exposure. However, the sheer size of the search space renders explicit optimisation —or "teaching to the test"—computationally infeasible. Given that real-world sensory inputs may encompass tens of thousands of bits, a genuine AGI system must therefore employ generalisable, pattern-based learning mechanisms capable of extracting latent structure from input data.
@@ 3.5 Performance
Before assessing the utility of AGITB, it is essential to recognise its role as a pragmatic benchmark rather than an end in itself. Similar to the Turing Test—which Harnad1992 described as an empirical criterion rather than a philosophical sleight of hand—AGITB is intended as a practical instrument for evaluating progress toward artificial general intelligence. The ultimate objective remains the development of AGI, not merely success on the benchmark.
AGITB can yield meaningful insights only when developers adhere strictly to its core requirements. Misinterpretations of fundamental elements, such as the "empty" initial state, may lead to erroneous conclusions and hinder genuine progress toward AGI.
Overall, AGITB provides a structured framework for empirically assessing the AGI potential of diverse computational paradigms, including classical symbolic architectures, artificial neural networks, and large language models. However, before benchmarking artificial systems, it is necessary to consider the baseline performance of human cognition.
@@ 3.5.1 Human Performance
The inability to directly compare internal cortical states makes it impossible to verify AGITB requirements in humans in a strict computational sense. Nevertheless, because cortical architecture inherently supports low-level binary signal processing and the tests align with cognitive expectations, it can be assumed that humans naturally satisfy most requirements. Demand 1 (Unbiased start), however, warrants further discussion.
Owing to prior experience and cognitive bias, an adult human may appear to "fail" the unbiased start prerequisite, as the cortex is no longer in an unconditioned state and may generate non-empty predictions. However, AGITB specifies that the unbiased state must precede the first input, a condition met in the fetal cortex. At that stage, the human cortex lacks synaptic organisation and, upon receiving its first sensory input, would satisfy the criterion of true neutrality by predicting no spikes.
Several of the more complex AGITB tests have cognitive-level analogues that can be observed through conscious reasoning and introspection. Temporal flexibility (Requirement 7), for example, is straightforward for humans, who can easily recognise temporal patterns of varying durations. Stagnation (Requirement 8) reflects the physical limitation of the human cortex to store and maintain knowledge. The corresponding test is analogous to the onset of cognitive saturation or early dementia, when recent experiences are forgotten. Because humans learn different topics at varying speeds, the rationale for content sensitivity (Requirement 9) and context sensitivity (Requirement 10) is self-evident. The unobservability test (Requirement 11) captures the fact that internal human mental states cannot be directly inspected. The denoising test (Requirement 12) and the generalisation test (Requirement 13) correspond to cognitive abilities in which humans excel—recalling and generalising when faced with new or distorted inputs. Finally, bounded prediction latency (Requirement 14) is a defining property of the human brain.
@@ 3.5.2 Classical Symbolic Programs
In principle, two complementary designs for AGI can be conceived. The first encodes, either explicitly or implicitly, certain expectations about the external world within the system itself. The second is entirely expectation-free, beginning from a neutral state without built-in assumptions or prior knowledge.
The former category encompasses most AI and putative AGI systems developed to date, yet it remains fundamentally constrained by the Symbol Grounding Problem (SGP), as articulated by Harnad1990. Although such systems may exhibit behaviour that appears intelligent, they lack grounded understanding and therefore cannot be regarded as genuine instances of AGI.
More specifically, systems that embed expectations inevitably reflect the programmer’s assumptions about the meaning of the signals from which they learn. In the case of standard symbolic programming, the program itself constitutes prior knowledge—its rules and representations presuppose an interpretation of the symbols being manipulated. The existence of such a program directly violates AGITB’s first test, which prohibits the inclusion of any external knowledge. Namely, the AGI program itself introduces the symbol-grounding problem it is meant to overcome.
AGITB’s initial test thus formalises the requirement to learn from scratch, demanding that a system derive structure and meaning solely through exposure to intrinsically grounded binary inputs. Only an expectation-free system could, in principle, satisfy this condition. To date, however, no such program is known. This suggests that a true AGI may not explicitly encode the operations of intelligence itself, but rather the dynamics of the substrate from which intelligence can emerge—a view consistent with the "Brain Simulator Reply" proposed by Churchland1990 in response to Searle’s Chinese Room Argument (Searle1980).
@@ 3.5.3 Artificial Neural Network Performance
A similar dichotomy applies to connectionist architectures. Artificial neural networks can, in principle, be constructed either with built-in expectations—through pretraining and architectural priors—or as expectation-free systems that begin from a neutral state. However, AGITB’s requirement that a model start in a completely unbiased, knowledge-free condition stands in fundamental tension with the prevailing paradigm of modern deep learning. Contemporary neural models typically rely on pretraining, whereby network weights are adjusted in advance through exposure to structured or labelled data.
Only an untrained network can be regarded as unbiased. However, an untrained network with all weights initialised to zero exhibits no learning dynamics: the neurons are not halted but remain dormant, producing no output. This exposes a fundamental limitation of current ANN architectures—they lack an intrinsic mechanism for initiating unbiased, autonomous learning from scratch in arbitrary and unfamiliar environments.
Furthermore, by translating symbols into numerical vectors, ANNs effectively transform the symbol grounding problem into a number grounding problem. While these numerical representations capture relational regularities among symbols, they also introduce spurious associations that are not anchored in real-world semantics, giving rise to the phenomenon commonly referred to as hallucination.
Pretraining on large datasets establishes correlations rather than genuine understanding. AGITB rejects this approach in favour of internal learning algorithms capable of initiating adaptation autonomously. However, no such mechanism has yet been demonstrated.
@@ 3.5.4 Large Language Model Performance
Because a large language model (LLM) is a type of artificial neural network (ANN), it fails the unbiased start test (Requirement 1). This failure demonstrates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained. Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—such evaluations remain moot until the first requirement is satisfied.
Two possibilities warrant consideration. The first concerns whether LLMs, while not qualifying as AGI, might nevertheless simulate one. For example, an LLM could in principle be prompted to imitate a model receiving binary inputs as specified by AGITB; a structured prompt of this type is provided in the Appendix. However, before evaluating such simulations, it is necessary to revisit the sensitivity requirement (Requirement 4), which stipulates that a model must remain responsive to any individual input even after processing an arbitrarily long sequence of subsequent inputs. This condition is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window. Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to lose or attenuate the long-term dependencies required by the test.
The second possibility is whether an LLM can autonomously generate an architecture capable of satisfying AGITB, given the reference implementation. In practice, this has not been observed: although systems such as ChatGPT-5 produced candidate programs purporting to pass the tests, none advanced beyond the adaptability test (Requirement 7).
To summarise, LLMs do not perform genuine learning solely from prompts, nor can they develop the grounded, context-dependent understanding of the world characteristic of human cognition. These limitations extend to large reasoning models (LRMs), which share the same underlying architectural principles.
@@ 3.6 Remarks
AGITB evaluates the model’s predictive capabilities following exposure to temporal sequences of either structured or random inputs. Random input sequences with arbitrary internal correlations are used to minimise reliance on pretraining, ensuring that any observed learning arises from the input stream rather than prior knowledge. By enforcing fundamental computational invariants of cortical function at the signal-processing level, AGITB remains agnostic to the external meaning of signals—the random inputs need not resemble real-world sensory data.
The low-level, binary operational framework makes AGITB particularly well-suited for evaluating NeuroAI models that aim to satisfy the principles of the embodied Turing Test, as proposed, for example, by Zador2023, where cognitive understanding emerges from the integration of continuous sensory streams. The progression from raw signal prediction to high-level abstraction reflects the broader evolution of AI itself—from early perceptrons to large-scale models such as GPT.
Hand-engineered, task-specific systems have historically struggled to scale toward true generalisation, reinforcing the need for adaptive, learning-based architectures such as artificial neural networks. However, state-of-the-art AI networks rely heavily on pretraining with structured or symbolic data—a process that inevitably introduces bias and circumvents the grounding of meaning in raw sensory input. This raises a critical and unresolved question: how can such models begin learning directly from unstructured input, as required to satisfy the AGITB?
@@ 3.7 Cheating the Benchmark
Since AGITB tests are simple to solve in isolation, a plausible strategy for circumventing the benchmark would be to develop task-specific solutions and make the model deploy them selectively based on the task at hand.
@@
While it is theoretically possible to infer the model’s assigned task and role by monitoring the number of instantiated model objects and their invoked methods, such an effort would serve only to subvert the benchmark rather than to advance AGI research. Although AGITB could incorporate additional obfuscation strategies—such as shuffling tests and redesigning the programming interface—these measures would not only deter cheating but also make tests less intelligible to human developers. This would undermine the testbed’s core objectives: transparency, readability, and interpretability.
@@
For this reason, the AGITB reference implementation remains deliberately readable and accessible. To date, no non-AGI system has demonstrated the level of performance that AGITB demands. Unless a credible attempt to circumvent the benchmark emerges, there is no justification for introducing an obfuscated version of the test.
4 Competing Benchmarks
Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC) introduced by Chollet2019 is most closely aligned in spirit, as it likewise emphasises generalisation over task-specific optimisation. A related effort is NeuroBench, which is designed to support the development of neuromorphic approaches to AGI. Both ARC and NeuroBench employ a variety of correctness and complexity metrics to differentiate among non-AGI systems, as they are primarily intended to distinguish weaker from stronger narrow models. In contrast, AGITB is designed to identify only true positive instances of AGI.
4.1 ARC
@@
ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recolouring, rearranging, or modifying spatial patterns in a spatial grid.
@@
However, ARC implicitly assumes the presence of high-level cognitive priors, such as object permanence, spatial reasoning, numerical abstraction, and causal inference. These priors are left undefined, placing an ambiguous burden on the model architecture. In contrast, AGITB adopts a fundamentally different perspective: it treats the model under evaluation as a blank slate that must develop structure and function exclusively through interaction with temporally structured input. Low-level reflexes are not regarded as prerequisites for intelligence within this framework, but rather as evolutionary features of the subcortical "old brain," as described by (Hawkins2004, p. 66).
@@
While ARC presumes temporal reasoning, it does not explicitly test it, as it uses only two images to demonstrate a transformation. In contrast, AGITB evaluates cognition as a dynamic, unfolding process over time. A model can only develop knowledge and predictive capabilities through time, not from disconnected before-and-after snapshots that lack the temporal continuity needed to infer causal structure. For instance, to learn about objects and to recognise the object transformation, such as "move left," a model under AGITB must observe multiple intermediate snapshots of the environment across time; the final image alone provides insufficient information to infer anything. Learning requires exposure to temporally structured data, enabling the model to discover invariants and causal relations through experience.
@@
In summary, ARC requires a high-level intelligence, whereas AGITB requires the 14 low-level principles. ARC evaluates performance based on behaviours grounded in human-like cognitive priors, while AGITB determines a system’s ability to support the emergence of these priors through exposure to raw input. Benchmarks such as ARC and the Turing test implicitly frame intelligence through the lens of human cognition, thereby embedding anthropocentric biases into their evaluation criteria. In contrast, AGITB conceptualises intelligence as a universal capacity for learning independent of innate symbolic structures or species-specific assumptions.
4.2 NeuroBench
NeuroBench provides a unified framework for benchmarking diverse AI models across a standardised set of tasks and metrics. It is particularly specialised for neuromorphic approaches, which have demonstrated advantages in resource efficiency and scalability. Within its algorithm track, the framework evaluates model performance on several ongoing challenges in general AI research, including few-shot continual learning, object detection, sensorimotor decoding, and predictive modelling.
The predictive modelling challenge, which involves forecasting chaotic functions, is most closely aligned with AGITB’s central premise that intelligence fundamentally concerns the prediction of future states. For this purpose, NeuroBench employs a synthetic one-dimensional Mackey–Glass time series—a dataset specifically designed for architectures with limited input/output capacity.
However, several issues limit NeuroBench’s suitability as a general AGI benchmark. First, the Mackey–Glass data are numerical; the meaning of a number, much like that of a symbol, is defined by human interpretation. The use of such representations necessarily reintroduces the symbol-grounding problem. Second, the threshold corresponding to AGI-level performance remains undefined. Although the symmetric mean absolute percentage error (sMAPE) is an established metric in forecasting and NeuroBench provides a baseline score, it does not specify what level of accuracy constitutes general intelligence. After all, humans perform poorly at detecting anomalies or making long-term predictions of the Mackey–Glass signal (Thill2020). Third, although long-term prediction is not problematic in itself, predicting multiple steps ahead without timely feedback deprives a system of the opportunity to recognise and correct its own errors. This design is incompatible with online learning, where an AGI should continuously update its internal state upon detecting discrepancies between its predictions and the actual outcomes. NeuroBench, by contrast, implicitly assumes that an AGI would behave as a purely mechanical predictor, lacking intrinsic mechanisms for self-correction, autonomous adaptation, or genuine agency.
4.3 Direct comparison
Table 1 summarises the principal distinctions among the three benchmarks. ARC and NeuroBench each presuppose the presence of high-level cognitive components—visual object recognition, manipulation and reasoning in the case of ARC, and comprehension and reasoning abilities in the case of NeuroBench. AGITB, by contrast, specifies elementary requirements that can be evaluated directly.
Property ARC CFP AGITB
Interface modality Visual Numeric Binary
Target reasoning type Human Universal Universal
Assumed cognitive priors Yes No No
Automation Limited Yes Yes
Teaching-to-the-test Yes No No
Intelligence level High Medium Low
Task preparation Manual Automatic Automatic
Symbol Grounding Problem Yes Yes No
Spatial dimension $200\times 100$ 1 10
Temporal dimension Yes Yes Yes
Table 1: A list of core properties of the ARC, NeuroBench’s chaotic function prediction (CFP), and AGITB.